OpenStudy (anonymous):

PRE-CALC/TRIG: Convert the polar equation to rectangular form and describe the curve: (@=theta) r=2sin@-4cos@

Directrix (directrix):

Circle. r = 2(y/r) -4(x/r) r^2 =2y -4x X^2 + y^2 = 2y -4x